Please help solve the equation, and check the solution for
3/8x - 1/3x +6= 1/6x

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

3/8x - 1/3x +6= 1/6x
------
Multiply thru by 24 to get:
9x - 8x + 144 = 4x
x + 144 = 3x
3x = 144
-----
x = 48
---------------------------
Cheers,
Stan H.
==================

Answer by rothauserc(4718)

(3/8)x - (1/3)x +6= (1/6)x
multiply both sides of the = by 24 and we get
9x - 8x + 144 = 4x
subtract x from both sides of = sign
3x = 144
x = 48
===========================================
check results
3/8 * 48 - 1/3 * 48 + 6 = 1/6 * 48
18 - 16 + 6 = 8
8 = 8
==========================================
x = 48 checks out :-)